I need help solving these system of linear equation problems.\

For this one I need to know if it has one solution or many.
3.) 2x + y - 3z = 4; 4x + 2z = 10; -2x + 3y - 13z = -8

Answer by richwmiller(17219)

3)
y = 19-8 x, z = 5-2 x
many solutions